OpenStudy (anonymous):

what is the seventh turn off the geometric sequence where a1 = -5 and a5 = -1,280?

OpenStudy (anonymous):

a5 = a1 * r^4 -1280 = -5 * r^4 solve for r a7 = a1 * r^6

OpenStudy (anonymous):

5120?

OpenStudy (anonymous):

It should be negative.

OpenStudy (anonymous):

i got it its -20,480!:)

OpenStudy (anonymous):

good.